# Stochastic Thermodynamics of Non-reciprocally Interacting Particles and Fields

### v2

@@ 4.2 Fluctuation response relation, higher order current cumulants and responses
In statistical physics, the FRR is a fundamental principle that connects equilibrium fluctuations with the linear response of the system. [175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 168]. The non-equilibrium analogue of FRR has been postulated [168, 185, 131, 132, 133, 134, 186, 187, 139, 188, 189, 190, 191, 192, 193, 194, 195, 196, 197]. We examine the FRR for the non-reciprocal and driven systems in this section. We define the $n^{th}$ scaled cumulant $\langle\prescript{n}{}{J}_{\gamma\gamma^{\prime}}\rangle_{C}$ of $J_{\gamma\gamma^{\prime}}$. By construction, $\langle\prescript{1}{}{J}_{\gamma\gamma^{\prime}}\rangle_{C}={J}_{\gamma\gamma^{\prime}}$ and $\langle\prescript{2}{}{J}_{\gamma\gamma^{\prime}}\rangle_{C}={T}_{\gamma\gamma^{\prime}}=2\Omega B_{\gamma\gamma^{\prime}}$. $\Omega$ defines the scaling between the scaled cumulant and the cumulant, for example, traffic ${T}_{\gamma\gamma^{\prime}}$ and variance $B_{\gamma\gamma^{\prime}}\left(\{\rho\}\right)$. It satisfies the following hierarchical relationship [86]:
$\begin{split}&\langle\prescript{n}{}{J}_{\gamma\gamma^{\prime}}\rangle_{C}=\langle\prescript{n-2}{}{J}_{\gamma\gamma^{\prime}}\rangle_{C},\\ &\langle\prescript{n}{}{J}_{\gamma\gamma^{\prime}}\rangle_{C}={J}_{\gamma\gamma^{\prime}},\hskip 28.45274pt\text{n is odd}.\\ &\langle\prescript{n}{}{J}_{\gamma\gamma^{\prime}}\rangle_{C}={T}_{\gamma\gamma^{\prime}},\hskip 28.45274pt\text{n is even}.\end{split}$ (44)
The eq. 44 reveals the recursive structure of the current cumulants. In particular, only the first and second cumulants are independent. This ensures the van Kampen moments closure expansion truncated up to the second order for the transition dynamics [92, 198, 199, 88]. Thus, eq. 44 highlights the validity of the van Kampen closure for far from equilibrium-driven non-reciprocal and externally driven systems. Physically, this important result enables the study of the macroscopic dynamics of non-reciprocal systems using the first two moments of the transition currents. A similar motif has previously been observed for the MFT [186]. This ensures the correctness of the Langevin/Gaussian approximation of the macrostate stochastic dynamics formulated in section B.
We define the response function $\prescript{A}{}{\zeta}_{\gamma\gamma^{\prime}}={\partial\>J_{\gamma\gamma^{\prime}}}/{\partial A_{\gamma\gamma^{\prime}}}$ and $\prescript{A}{}{\vec{\zeta}}_{i}^{\vec{\mathcal{D}}}={\partial\>\vec{J}_{i}^{\vec{\mathcal{D}}}}/{\partial\vec{A}_{i}^{\vec{\mathcal{D}}}}$. The response function satisfies FRR:
@@
The $\Omega$ plays an analogous role to the inverse temperature for the macroscopic stochastic dynamics [167]. We use the convention of evaluating the response at the reference affinity $A_{\gamma\gamma^{\prime}}^{*}$. Hence, it characterizes the reference state and the probability distribution around which the response is evaluated, for instance, a steady state or an equilibrium distribution. We define the response function for the currents with respect to the change in symmetric reactive transition affinity, $\prescript{U}{}{\zeta}_{\gamma\gamma^{\prime}}={\partial\>J_{\gamma\gamma^{\prime}}}/{\partial U_{\gamma\gamma^{\prime}}}$ and $\prescript{U}{}{\vec{\zeta}}_{i}^{\vec{\mathcal{D}}}={\partial\>\vec{J}_{i}^{\vec{\mathcal{D}}}}/{\partial\vec{U}_{i}^{\vec{\mathcal{D}}}}$.
@@
The set of eqs. 45 and 46 satisfies the generic linear response relations between current and traffic [130, 195].
We further delineate the underlying generic structure for the higher-order response function. The $n^{th}$ order response function for the $m^{th}$ current cumulants is defined as $\prescript{A}{n}{\zeta}_{\gamma\gamma^{\prime}}^{m}={\partial^{n}\>\langle\prescript{m}{}{J}_{\gamma\gamma^{\prime}}\rangle_{C}}/{\partial A_{\gamma\gamma^{\prime}}^{n}}$ and $\prescript{U}{n}{\zeta}_{\gamma\gamma^{\prime}}^{m}={\partial^{n}\>\langle\prescript{m}{}{J}_{\gamma\gamma^{\prime}}\rangle_{C}}/{\partial U_{\gamma\gamma^{\prime}}^{n}}$,
@@
eq. 47 is the higher-order FRR that relates the non-linear response of any higher-order current cumulant to other current cumulants.
where $\delta(n_{1}-n_{2})$ is the Kronecker delta function. The eqs. 48 and 47 generalizes the far-from-equilibrium higher-order FRR for non-reciprocal and driven systems [200, 201, 202].
We use the FRR to infer the transition affinity $A_{\Delta}^{*}$ using the mean current $\langle J_{\Delta}\rangle$ and current variance $2B_{\Delta}$ of the transition $\Delta$ [88].
Using eq. 49, we formulate an inference-based inverse problem. The affinity $A^{*}_{\Delta^{o}}$ of an observable transition $\Delta_{0}$ is inferred using the observable mean current $J_{\Delta^{o}}$ and it’s variance $2B_{\Delta^{o}}$ [88],
$A_{\Delta^{o}}^{*}$ is the effective driving force corresponding to the observable current $J_{\Delta^{o}}^{*}$. This concludes the formulation of the FRR for the macroscopic fluctuating dynamics of non-reciprocal systems.
4.3 Fluctuation relations
The fluctuation relations (FR) stand as fundamental principles that illuminate the behaviour of far-from-equilibrium fluctuating systems [2, 3, 4, 5, 6, 7, 15, 8, 11, 10, 203, 204]. It offers a profound insight into the nature of fluctuations, shedding light on the asymmetry (symmetry) between the forward and backward physical processes at the microscopic level [12, 11, 10, 14, 13, 15]. By examining the statistical properties of systems undergoing non-equillibrium stochastic dynamics, FR unveils universal laws about the irreversible nature of thermodynamic processes. This bridges the gap between macroscopic irreversibility and the underlying microscopic dynamics, paving the way for a deeper understanding of the profound interplay between order and fluctuations in the physical world. The FR generalizes the FRR and Onsager’s regression hypothesis for systems operating far from equilibrium [205, 206, 207, 208, 209].
Here, we will focus on a unified formalism of FR based on the measure theory [210, 211, 212, 213] combined with the large deviation principle [124]. In measure theory, the Radon-Nikodym derivative (RND) is defined as the transition probability measure between the process and the corresponding reference process. The LDB condition establishes the connection between the RND as a mathematical property: a transition probability measure and the physical interpretation of it as a stochastic transition EP. The measure-theoretical formalism of the FR based on RND utilizes the contraction of the rate functional for the empirical microscopic transition currents to the corresponding rate functional for the EP [88]. The orthogonal decomposition of the stochastic EP ensures that the operation and the reference operation chosen to evaluate RND are spanned by the operations that commute with each other. From the large deviation approach, it is equivalent to choosing the linearly independent empirical observable EP [214, 215, 216, 217, 218]. The orthogonality condition delineates the linearly independent symmetry operations that act on the system [214, 215, 216, 217, 218].
[…]
$\begin{split}G_{\vec{{{O}}}}(\vec{\chi}_{{O}})=\ln{\langle e^{\vec{\chi}_{{O}}\>\cdot\>\vec{\Omega}\>\odot\>\vec{\hat{O}}}\rangle},\end{split}$ (51)
with the observable conjugate vector $\vec{\chi}_{\vec{{O}}}=\{\chi_{\Delta\hat{S}^{state}_{E}},\chi_{\Delta\hat{\Psi}_{E}},\chi_{\hat{W}},\{\chi_{\hat{\omega}_{ij}}\},\{\chi_{\hat{J}_{\gamma\gamma^{\prime}}}\},\{\chi_{\hat{J}_{i}^{\vec{\mathcal{D}}}}\}\}$. The scaling of $G_{\vec{{{O}}}}(\vec{\chi}_{{O}})$ is employed using $max\{\vec{\Omega}_{O}\}$ that dominates the Scaled $G_{\vec{{{O}}}}(\vec{\chi}_{{O}})$ (SCGF). The probability density measure $\mathcal{P}(\vec{O})$ satisfies the following symmetry:
@@ 4.3 Fluctuation relations
Defined precisely $\mathcal{P}(\vec{O})=\lim_{\delta\vec{O}\to 0}\mathcal{P}(\vec{O}<\vec{\hat{O}}<\vec{O}+\delta\vec{O})$ and $\odot$ is the component-wise Haddamard product defined between two vectors. The eq. 52 is the finite $\vec{\Omega}_{O}$ Gallovotti-Cohen fluctuation relation symmetry, which requires identifying $\vec{A}_{\vec{\hat{O}}}=\{1,-1,1,\{V^{nr}_{ij}\},\{F^{ch}_{\gamma\gamma^{\prime}}\},\{\vec{F}^{sp}_{i}\}\}$, the forces conjugate to the $\vec{\hat{O}}$ [12, 207, 213, 219, 220, 221, 208, 222, 223, 224, 201, 225, 226, 202].
We aim to identify the effective macroscopic symmetry of coarse-grained observable denoted by $\vec{O}^{\parallel}$, such that $\vec{O}$ is decomposed into $\vec{O}=\vec{O}^{\parallel}\oplus\vec{O}^{\perp}$. Using the Bayes theorem and eq. 52, the FR for $\vec{O}^{\parallel}$ reads [214, 215, 216, 217, 218]:
@@
where, $\langle e^{-\vec{O}^{\perp}}|\vec{O}^{\parallel}\rangle=\int\mathbb{D}\vec{O}^{\perp}\>{\mathcal{P}(\vec{O}^{\perp}|\vec{O}^{\parallel})}e^{-\vec{A}_{\vec{\hat{O}}^{\perp}}\cdot\vec{\Omega}^{\perp}\odot\vec{O}^{\perp}}$ quantifies the projection of the conditional probability measure of $\vec{O}^{\perp}$ into the $\vec{O}^{\parallel}$ space. $\langle e^{-\vec{O}^{\perp}}|\vec{O}^{\parallel}\rangle=1$ for the linearly independent observable $\vec{O}^{\parallel}$ and $\vec{O}^{\perp}$, which recovers the FR for $\vec{O}^{\parallel}$. The proof follows trivially by the Taylor series expansion of $e^{-\vec{O}^{\perp}}$ and orthogonality relation $\langle\vec{O}^{\perp}|\vec{O}^{\parallel}\rangle=0$. This emphasizes the importance of choosing the linearly independent orthogonal observable. Physically, this generates a hierarchy of detailed fluctuation relations and delineates independent underlying symmetries of EP.
By choosing, $\vec{O}^{\parallel}=\vec{\hat{\Sigma}}=\{\hat{W},\Delta_{0}^{\tau}\hat{S}^{state}_{E},\hat{\Sigma}^{nr},\hat{\Sigma}^{ch},\hat{\Sigma}^{sp}\}$ and $\vec{O}^{\perp}=\Delta_{0}^{\tau}\psi_{E}$, we obtain the detailed master FR for non-reciprocal systems:
@@
We have utilized the contraction from the linearly independent empirical current observable eq. 52 to the linearly independent EP eq. 54 [88]. The eq. 54 satisfies the Lebowitz-Spohn symmetry, with the generating function satisfying $G_{\vec{\Sigma}}(\vec{\chi}_{\vec{\Sigma}})=G_{\vec{{\Sigma}}}(-1-\vec{\chi}_{\vec{{\Sigma}}})$ [13, 15, 11]. The stochastic work $W$ is not a linearly independent variable, rather the irreversible work $W-\Delta_{0}^{\tau}\psi_{E}$ is [217, 218]. This is effectively realized by the factor of $W-\Delta_{0}^{\tau}\psi_{E}$ on the left-side of the eq. 54. Hence $\Delta_{0}^{\tau}\psi_{E}$ drops out and its projection onto the work gives the $W-\Delta_{0}^{\tau}\psi_{E}$ on the right-side of eq. 54.
Considering $F_{\gamma\gamma^{\prime}}^{ch}=0$ and $\vec{F}_{i}^{sp}=0$, the eq. 54 is reduced to the Crooks-Tasaki relation for the non-reciprocal systems [9, 10, 8]:
@@
Assuming the external driving control parameter change instantaneously relaxes the probability distribution to the reference Boltzmann distribution (no error during driving or a quasistatic driving process [217, 227, 218]), $\Delta_{0}^{\tau}S^{state}_{E}=0$. The eq. 55 is reduced to a familiar form of the Crooks-Tasaki fluctuation theorem. By integrating eq. 55, one obtains the Jarzynski fluctuation relation for non-reciprocal systems:
@@
Using eq. 42, we relate the external driving work needed with the symmetric and anti-symmetric correlations between the initial and final macrostates. Using Jensen’s inequality, the second law of thermodynamics (approximate law) eq. 40 is recovered. However, using $C_{ij}^{a}(t,\tau)$ and the fluctuation theorem (exact law) eq. 56 gives a tighter bound on the mean work. The work done on non-reciprocal systems requires extra thermodynamic cost associated with changing the anti-symmetric temporal cross-correlations between the macrostate  [^3] The validity of the non-reciprocal Crooks-Tasaki and Jarzynski relation eqs. 55 and 56 is subjected to not crossing the $\mathcal{PT}$ symmetry braking bifurcation manifold in control parameters space. A driving through $\mathcal{PT}$ transition from a static to dynamic phase undergoes diverging fluctuations which could lead to nontrivial FR symmetries and exponents. A more systematic study is needed for such physical problems.. By choosing $\Delta_{0}^{\tau}S^{state}_{E}\neq 0$, the stochastic work incorporates the irreversible work contribution due to the non-quasistatic driving of control parameters [79, 229, 80, 230, 227, 218, 231, 217]. Physically, it quantifies the mismatch between the instantaneous non-equilibrium probability distribution and the corresponding instantaneous Boltzmann probability distribution specified by the instantaneous control parameters of $E$. In the absence of the external work $(W=0)$, $\Delta_{0}^{\tau}S^{state}_{E}\neq 0$ quantifies the relaxation EP and restores the finite-time FR.
Considering $W=0$ and combining the boundary term $\Delta\psi_{E}$ with the non-conservative EP $\Sigma^{nr}+\Sigma^{sp}+\Sigma^{ch}$, eq. 54 is reduced to the following equation:
@@
The eq. 57 is a generic three detailed fluctuation relations for the non-adiabatic, housekeeping and total EP applicable to the non-reciprocal systems [102]. The eq. 57 is the effective version of the underlying FR symmetry eq. 54. The detailed FR for the non-adiabatic, housekeeping and total EP are obtained by using $E=ss$ [229, 232, 102] or Hatana-Sasa relation for a transition from an initial to final steady state [233, 102]. The steady-state condition is not required to formulate the orthogonal decomposition, making the orthogonal decomposition applicable in a broader context, with a suitable gauge fixing choice of $E$ chosen according to the required physical/experimental constraint [104, 103, 105, 107, 108, 109, 106, 148, 149, 150, 151, 152, 153, 154, 155, 102].

### v3

@@ 4.2 Fluctuation response relation, higher order current cumulants and responses
In statistical physics, the FRR is a fundamental principle that connects equilibrium fluctuations with the linear response of the system [181, 182, 183, 184, 185, 186, 187, 188, 189, 190, 174]. The non-equilibrium analog of FRR has been postulated [174, 191, 137, 138, 139, 140, 192, 193, 145, 194, 195, 196, 197, 198, 199, 200, 201, 202, 203]. We examine the FRR for non-reciprocal and driven systems in this section. We define the $n^{th}$ scaled cumulant $\langle\prescript{n}{}{J}_{\gamma\gamma^{\prime}}\rangle_{C}$ of $J_{\gamma\gamma^{\prime}}$. By construction, $\langle\prescript{1}{}{J}_{\gamma\gamma^{\prime}}\rangle_{C}={J}_{\gamma\gamma^{\prime}}$ and $\langle\prescript{2}{}{J}_{\gamma\gamma^{\prime}}\rangle_{C}={T}_{\gamma\gamma^{\prime}}=2\Omega B_{\gamma\gamma^{\prime}}$. Here, $\Omega$ defines the scaling between the scaled cumulant and the cumulant, for example, the traffic ${T}_{\gamma\gamma^{\prime}}$ and the variance $B_{\gamma\gamma^{\prime}}\left(\{\rho\}\right)$. It satisfies the following hierarchical relationship [87]:
$\begin{split}&\langle\prescript{n}{}{J}_{\gamma\gamma^{\prime}}\rangle_{C}=\langle\prescript{n-2}{}{J}_{\gamma\gamma^{\prime}}\rangle_{C},\\ &\langle\prescript{n}{}{J}_{\gamma\gamma^{\prime}}\rangle_{C}={J}_{\gamma\gamma^{\prime}},\hskip 28.45274pt\text{n is odd},\\ &\langle\prescript{n}{}{J}_{\gamma\gamma^{\prime}}\rangle_{C}={T}_{\gamma\gamma^{\prime}},\hskip 28.45274pt\text{n is even}.\end{split}$ (45)
Equation 45 reveals the recursive structure of the current cumulants. In particular, only the first and second cumulants are independent. This ensures that the van Kampen moments closure expansion is truncated up to the second order for the transition dynamics [97, 204, 205, 89, 90]. Thus, eq. 45 highlights the validity of the van Kampen closure for far-from-equilibrium, non-reciprocal, and externally driven systems. Physically, this important result enables the study of the macroscopic dynamics of non-reciprocal systems using the first two moments of the transition currents. A similar motif has previously been observed for MFT [192]. This ensures the correctness of the Langevin/Gaussian approximation of the macrostate stochastic dynamics formulated in the dynamics section.
We define the response function $\prescript{A}{}{\zeta}_{\gamma\gamma^{\prime}}={\partial\>J_{\gamma\gamma^{\prime}}}/{\partial A_{\gamma\gamma^{\prime}}}$ and $\prescript{A}{}{\vec{\zeta}}_{i}^{\vec{\mathcal{D}}}={\partial\>\vec{J}_{i}^{\vec{\mathcal{D}}}}/{\partial\vec{A}_{i}^{\vec{\mathcal{D}}}}$. The response function satisfies the FRR:
@@
Here, $\Omega$ plays an analogous role to the inverse temperature for the macroscopic stochastic dynamics [173]. We use the convention of evaluating the response at the reference affinity $A_{\gamma\gamma^{\prime}}^{*}$. Hence, it characterizes the reference state and the probability distribution around which the response is evaluated, for instance, a steady state or an equilibrium distribution. We define the response function for the currents with respect to the change in symmetric reactive transition affinity, $\prescript{U}{}{\zeta}_{\gamma\gamma^{\prime}}={\partial\>J_{\gamma\gamma^{\prime}}}/{\partial U_{\gamma\gamma^{\prime}}}$ and $\prescript{U}{}{\vec{\zeta}}_{i}^{\vec{\mathcal{D}}}={\partial\>\vec{J}_{i}^{\vec{\mathcal{D}}}}/{\partial\vec{U}_{i}^{\vec{\mathcal{D}}}}$:
@@
The set of eqs. 46 and 47 satisfies the generic linear response relations between current and traffic [136, 201].
We further delineate the underlying generic structure for the higher-order response function. The $n^{th}$ order response function for the $m^{th}$ current cumulant is defined as $\prescript{A}{n}{\zeta}_{\gamma\gamma^{\prime}}^{m}={\partial^{n}\>\langle\prescript{m}{}{J}_{\gamma\gamma^{\prime}}\rangle_{C}}/{\partial A_{\gamma\gamma^{\prime}}^{n}}$ and $\prescript{U}{n}{\zeta}_{\gamma\gamma^{\prime}}^{m}={\partial^{n}\>\langle\prescript{m}{}{J}_{\gamma\gamma^{\prime}}\rangle_{C}}/{\partial U_{\gamma\gamma^{\prime}}^{n}}$:
$\begin{split}\prescript{A}{n}{\zeta}_{\gamma\gamma^{\prime}}^{m}=\frac{1}{2^{n}}\langle\prescript{m+n}{}{J}_{\gamma\gamma^{\prime}}\rangle_{C},\hskip 22.76228pt\prescript{U}{n}{\zeta}_{\gamma\gamma^{\prime}}^{m}=\frac{1}{2^{n}}\langle\prescript{m}{}{J}_{\gamma\gamma^{\prime}}\rangle_{C},\end{split}$ (48)
eq. 48 is the higher-order FRR that relates the non-linear response of any higher-order current cumulant to other current cumulants:
where $\delta(n_{1}-n_{2})$ is the Kronecker delta function. Equations 49 and 48 generalizes the far-from-equilibrium higher-order FRR for non-reciprocal and driven systems [206, 207, 208].
We use the FRR to infer the transition affinity $A_{\Delta}^{*}$ using the mean current $\langle J_{\Delta}\rangle$ and the current variance $2B_{\Delta}$ of the transition $\Delta$ [89, 90]:
Using eq. 50, we formulate an inference-based inverse problem. The affinity $A^{*}_{\Delta^{o}}$ of an observable transition $\Delta_{0}$ is inferred using the observable mean current $J_{\Delta^{o}}$ and its variance $2B_{\Delta^{o}}$ [89, 90]:
$A_{\Delta^{o}}^{*}$ is the effective driving force corresponding to the observable current $J_{\Delta^{o}}$. This concludes the formulation of the FRR for the macroscopic fluctuating dynamics of non-reciprocal systems.
4.3 Fluctuation relations
The fluctuation relations (FR) stand as fundamental principles that illuminate the behavior of far-from-equilibrium fluctuating systems [2, 3, 4, 5, 7, 8, 11, 12, 13, 16, 209, 210]. They offer profound insight into the nature of fluctuations, shedding light on the asymmetry (symmetry) between the forward and backward physical processes at the microscopic level [6, 13, 16, 17, 14, 11]. By examining the statistical properties of systems undergoing non-equilibrium stochastic dynamics, the FR unveil universal laws about the irreversible nature of thermodynamic processes. This bridges the gap between macroscopic irreversibility and the underlying microscopic dynamics, paving the way for a deeper understanding of the interplay between order and fluctuations in physical systems. The FR generalize the fluctuation-dissipation relation (FRR) and Onsager’s regression hypothesis for systems operating far from equilibrium [211, 212, 213, 214, 215].
Here, we focus on a unified formalism of the FR based on measure theory [216, 217, 218, 219] combined with the large deviation principle [130]. In measure theory, the Radon-Nikodym derivative (RND) is defined as the transition probability measure between the process and the corresponding reference process. The LDB condition establishes the connection between the RND as a mathematical property, that is, a transition probability measure, and its physical interpretation as a stochastic transition EP. The measure-theoretical formalism of the FR based on the RND utilizes the contraction of the rate functional for the empirical microscopic transition currents to the corresponding rate functional for the EP [89, 90]. The orthogonal decomposition of the stochastic EP ensures that the operation and the reference operation chosen to evaluate the RND are spanned by operations that commute with each other. From the large deviation approach, this is equivalent to choosing the linearly independent empirical observable EP [220, 221, 222, 223, 224, 90]. The orthogonality condition delineates the linearly independent symmetry operations that act on the system [220, 221, 222, 223, 224, 90].
We define a scaled-intensive observable $\hat{O}=\frac{1}{\Omega_{O}}\int_{0}^{\tau}Odt$ with a scaling factor $\Omega_{O}$. We consider the most general observable $\vec{\hat{O}}=\{\Delta_{0}^{\tau}\hat{S}^{state}_{E},\Delta_{i}^{\tau}\psi_{E},\hat{W},\{\hat{\omega}_{ij}\},\{\hat{J}_{\gamma\gamma^{\prime}}^{ch}\},\{\vec{\hat{J}}_{i}^{sp}\}\}$ in vector representation with $\vec{\Omega}_{O}=\{\ln{(\Omega)},\Omega,\Omega,\{\Omega\tau\},\{\Omega\tau\},\{\Omega\tau\}\}$ and $\vec{\Omega}_{O}^{-1}=\{(\ln{(\Omega}))^{-1},\Omega^{-1},\Omega^{-1},(\Omega\tau)^{-1},(\Omega\tau)^{-1},(\Omega\tau)^{-1}\}$. Here, we have utilized the counting observable for the transition currents. The cumulant generating function $G_{\vec{{\mathbf{O}}}}(\vec{\chi}_{{O}})$ for $\vec{{O}}$ reads:
@@ 4.3 Fluctuation relations
with the observable conjugate vector $\vec{\chi}_{\vec{{O}}}=\{\chi_{\Delta\hat{S}^{state}_{E}},\chi_{\Delta\hat{\Psi}_{E}},\chi_{\hat{W}},\{\chi_{\hat{\omega}_{ij}}\},\{\chi_{\hat{J}_{\gamma\gamma^{\prime}}}\},\{\chi_{\hat{J}_{i}^{\vec{\mathcal{D}}}}\}\}$. The scaling of $G_{\vec{{{O}}}}(\vec{\chi}_{{O}})$ is employed using $\max\{\vec{\Omega}_{O}\}$, which dominates the scaled $G_{\vec{{{O}}}}(\vec{\chi}_{{O}})$ (SCGF). The probability density measure $\mathcal{P}(\vec{O})$ satisfies the following symmetry:
@@
Defined precisely, $\mathcal{P}(\vec{O})=\lim_{\delta\vec{O}\to 0}\mathcal{P}(\vec{O}<\vec{\hat{O}}<\vec{O}+\delta\vec{O})$ and $\odot$ is the component-wise Hadamard product defined between two vectors. Equation 53 is the finite-$\vec{\Omega}_{O}$ Gallavotti-Cohen fluctuation relation symmetry, which requires identifying $\vec{A}_{\vec{\hat{O}}}=\{1,-1,1,\{V^{nr}_{ij}\},\{F^{ch}_{\gamma\gamma^{\prime}}\},\{\vec{F}^{sp}_{i}\}\}$, the forces conjugate to $\vec{\hat{O}}$ [6, 213, 219, 225, 226, 227, 214, 228, 229, 230, 207, 231, 232, 208].
We aim to identify the effective macroscopic symmetry of the coarse-grained observable denoted by $\vec{O}^{\parallel}$, such that $\vec{O}$ is decomposed into $\vec{O}=\vec{O}^{\parallel}\oplus\vec{O}^{\perp}$. Using Bayes’ theorem and eq. 53, the FR for $\vec{O}^{\parallel}$ reads [220, 221, 222, 223, 224]:
@@
where $\langle e^{-\vec{O}^{\perp}}|\vec{O}^{\parallel}\rangle=\int\mathbb{D}\vec{O}^{\perp}\>{\mathcal{P}(\vec{O}^{\perp}|\vec{O}^{\parallel})}e^{-\vec{A}_{\vec{\hat{O}}^{\perp}}\cdot\vec{\Omega}^{\perp}\odot\vec{O}^{\perp}}$ quantifies the projection of the conditional probability measure of $\vec{O}^{\perp}$ into the $\vec{O}^{\parallel}$ space. $\langle e^{-\vec{O}^{\perp}}|\vec{O}^{\parallel}\rangle=1$ for linearly independent observables $\vec{O}^{\parallel}$ and $\vec{O}^{\perp}$, which recovers the FR for $\vec{O}^{\parallel}$. The proof follows trivially by the Taylor series expansion of $e^{-\vec{O}^{\perp}}$ and the orthogonality relation $\langle\vec{O}^{\perp}|\vec{O}^{\parallel}\rangle=0$. This emphasizes the importance of choosing linearly independent orthogonal observables. Physically, this generates a hierarchy of detailed fluctuation relations and delineates independent underlying symmetries of the EP.
By choosing $\vec{O}^{\parallel}=\vec{\hat{\Sigma}}=\{\hat{W},\Delta_{0}^{\tau}\hat{S}^{state}_{E},\hat{\Sigma}^{nr},\hat{\Sigma}^{ch},\hat{\Sigma}^{sp}\}$ and $\vec{O}^{\perp}=\Delta_{0}^{\tau}\psi_{E}$, we obtain the detailed master FR for non-reciprocal systems:
@@
We have implemented the contraction from the linearly independent empirical current observable eq. 53 to the linearly independent EP eq. 55 [89, 90]. Equation 55 satisfies the Lebowitz-Spohn symmetry, with the generating function satisfying $G_{\vec{\Sigma}}(\vec{\chi}_{\vec{\Sigma}})=G_{\vec{{\Sigma}}}(-1-\vec{\chi}_{\vec{{\Sigma}}})$ [14, 11, 13]. The stochastic work $W$ is not a linearly independent variable; rather, the irreversible work $W-\Delta_{0}^{\tau}\psi_{E}$ is [223, 224]. This is effectively realized by the factor of $W-\Delta_{0}^{\tau}\psi_{E}$ on the left-hand side of eq. 55. Hence, $\Delta_{0}^{\tau}\psi_{E}$ drops out, and its projection onto the work gives the factor $W-\Delta_{0}^{\tau}\psi_{E}$ on the right-hand side of eq. 55.
Assuming the specific case of $F_{\gamma\gamma^{\prime}}^{ch}=0$ and $\vec{F}_{i}^{sp}=0$, which corresponds to non-reciprocity being the only source of the system’s non-equilibrium-ness, eq. 55 is reduced to the Crooks-Tasaki relation for non-reciprocal systems [15, 16, 12]:
@@
Assuming that the external driving control parameter changes instantly relax the probability distribution to the reference Boltzmann distribution (no error during driving, i.e., a quasistatic driving process [223, 233, 224]), we have $\Delta_{0}^{\tau}S^{state}_{E}=0$. Equation 56 is then reduced to a familiar form of the Crooks-Tasaki fluctuation theorem, but defined here for non-reciprocal systems. By integrating eq. 56, we obtain the Jarzynski fluctuation relation for non-reciprocal systems:
@@
Using eq. 43, we relate the external driving work needed with the symmetric and anti-symmetric correlations between the initial and final macrostates. Using Jensen’s inequality, the second law of thermodynamics (approximate law) eq. 41 is recovered. However, using $C_{ij}^{a}(t,\tau)$ and the fluctuation theorem (exact law) eq. 57 gives a tighter bound on the mean work. The work done on non-reciprocal systems requires extra thermodynamic cost associated with changing the anti-symmetric temporal cross-correlations between the macrostates. The validity of the non-reciprocal Crooks-Tasaki and Jarzynski relations eqs. 56 and 57 is subject to not crossing the $\mathcal{PT}$ symmetry-breaking bifurcation manifold in the space of control parameters. Driving through a $\mathcal{PT}$ transition from a static to dynamic phase undergoes divergent fluctuations, which could lead to non-trivial FR symmetries and exponents. A more systematic study is needed for such physical problems.
By choosing $\Delta_{0}^{\tau}S^{state}_{E}\neq 0$, the stochastic work incorporates the irreversible work contribution due to the non-quasistatic driving of the control parameters [80, 234, 81, 235, 233, 224, 236, 223]. Physically, it quantifies the mismatch between the instantaneous non-equilibrium probability distribution and the corresponding instantaneous Boltzmann probability distribution specified by the instantaneous control parameters of $E$. In the absence of external work $(W=0)$, $\Delta_{0}^{\tau}S^{state}_{E}\neq 0$ quantifies the relaxation EP and restores the finite-time FR.
Assuming $W=0$ and combining the boundary term $\Delta\psi_{E}$ with the non-conservative EP $\Sigma^{nr}+\Sigma^{sp}+\Sigma^{ch}$, eq. 55 is reduced to the following equation:
@@
Equation 58 represents generic detailed fluctuation relations for the non-adiabatic, housekeeping, and total EP, applicable to non-reciprocal systems [108]. It is the effective version of the underlying FR symmetry eq. 55. The detailed FR for the non-adiabatic, housekeeping, and total EP are obtained by using $E=ss$ [234, 237, 108], the Hatano-Sasa relation for the excess EP for a transition from an initial to a final steady state [238, 108], or the Speck-Seifert relation for the housekeeping EP for a transition from an initial to a final steady state [239].
The steady-state condition is not required to formulate the orthogonal decomposition, making it applicable in a broader context, with a suitable gauge-fixing choice of $E$ according to the required physical or experimental constraints [90, 109, 111, 113, 114, 115, 112, 154, 155, 156, 157, 158, 159, 160, 161, 108].